A CAR STARTS RECTILINEARLY , FIRST WITH ACC. 'W' = 5 M/S2 (THE INTIAL VELOCITY  IS 0), THEN UNIFORMLY,ND DEN DECELARATING AT D SAME RATE COMES TO A STOP. THE TOTAL TIME MOTION EQUALS T = 25S. THE AV. VELOCITY IS 72 KM/H. HOW LONG DOES D CAR MOVE UNFORMLY??


 


THE PROBLEM IS WHETHER THE ACCELERATION IS UNIFORM OR THE VELOCITY DAT IS WAT IM NOT ABLE TO UNDERSTAND??

Grade:12

1 Answers

Prudhvi teja
13 years ago

Dear shashwat

He means you need to find out till what time it went through constant speed.

Take time with constant velocity is t

and with acceleration is t1

so 2t1+t= 25

avg velocity = ((at1)t1+(at1)t)/(2t1+t) = 20m/s

t1(t1+t) = 100

t1(25-t1) = 100

t1=5

t= 15s

Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly.

All the best.

Regards,
Askiitians Experts
Prudhvi Teja

 

 


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian